# Coherent Load Profile Synthesis with Conditional Diffusion for LV Distribution Network Scenario Generation

### v1

@@ Abstract
Limited visibility of power distribution network power flows at the low voltage level presents challenges to both distribution network operators from a planning perspective and distribution system operators from a congestion management perspective. Forestalling these challenges through scenario analysis is confounded by the lack of realistic and coherent load data across representative distribution feeders. Load profiling approaches often rely on summarising demand through typical profiles, which oversimplifies the complexity of substation-level operations and limits their applicability in specific power system studies. Sampling methods, and more recently generative models, have attempted to address this through synthesising representative loads from historical exemplars; however, while these approaches can approximate load shapes to a convincing degree of fidelity, the co-behaviour between substations, which ultimately impacts higher voltage level network operation, is often overlooked. This limitation will become even more pronounced with the increasing integration of low-carbon technologies, as estimates of base loads fail to capture load diversity. To address this gap, a Conditional Diffusion model for synthesising daily active and reactive power profiles at the low voltage distribution substation level is proposed. The evaluation of fidelity is demonstrated through conventional metrics capturing temporal and statistical realism, as well as power flow modelling. The results show synthesised load profiles are plausible both independently and as a cohort in a wider power systems context. The Conditional Diffusion model is benchmarked against both naive and state-of-the-art models to demonstrate its effectiveness in producing realistic scenarios on which to base sub-regional power distribution network planning and operations.
@@
[…]
[1]organization=National Manufacturing Institute Scotland, city=Renfrew, postcode=PA3 2EF, country=Scotland
@@ 1 Introduction
Scenario generation is one viable route to forestall potential problems from the introduction of new technology or general load growth. However, a baseline performance measure is difficult to obtain due to the lack of LV observability. At the distribution level, the primary concern is the extremes: headroom for thermal exceedances and footroom for voltage violations [5]. Typical Load Profiles (TLP) exist [6], but these are generally seasonal, not weather sensitive and lack the diversity that would enable a cohort of premises to be aggregated together into a plausible LV substation load. In defining plausibility, the peak demand sharpness (and under embedded PV scenarios, trough depth) should reflect the range of time use across neighbourhood premises: too sharp would result from near identical behaviours and hence an overestimate of potential threats; too blunt assumes greater diversity than is likely, and therefore a potential overestimate of headroom. Generative modelling from a limited set of historical exemplars offers a potential solution, but the varied nature of LV substation loading introduces its own set of complexities and hurdles. The statistical distribution of LV substation load does not follow any common probability density function [7], making sampling approaches unrepresentative unless carefully chosen for particular locations. Markov Chain approaches also struggle with capturing aggregated time use resulting from routine behaviours [8]. Understanding the present loading characteristics is essential as this condition could then further worsen with the extensive installation of distributed generators as well as low-carbon heat and transport [9]. This will cause equipment to operate under atypical conditions that impact the wider network [10], potentially up to the transmission level. For a Distribution Network Operator (DNO), this insight is necessary for long-term planning of reinforcements that will facilitate LCT adoption and generally increase resilience over time periods of months, years and decades. For a Distribution System Operator (DSO), the same insights will yield short-term generation dispatch and load curtailment potential within the constraints of the physical network infrastructure, with both active and reactive turn-up and turn-down capabilities articulated accurately via a flexibility envelope  [11]. This, in turn, will articulate substation flexibilities that can provide balancing at the transmission interface, which, combined, represent regional-scale scenarios from fundamental end-use insights - unlocking this knowledge of aggregating demand would enable better system planning as well as policy design over a longer period.
@@
LV distribution network loads are much closer to individual premises end use, meaning that lower levels of aggregation will reflect highly variable behaviour routines which may result from social norms, variation in appliance ownership and specification, and also reactions to localised weather conditions. Previously, LV substation data was modelled using conventional methods such as aggregated average profiling [6] with a diversity correction. While this approach was historically sufficient, the increasing penetration of LCTs now necessitates more advanced and accurate methods. To approach this complex and non-stationary behaviour, it is proposed here that a generative AI model is applied, which is capable of synthesising both Active and Reactive power data for LV substation load data based on a set of observed influencing factors, namely, weather and calendar variables. This is novel in the following ways: Firstly, applying a deep generative model to LV substation load modelling. More specifically, a Diffusion model, which is the state-of-the-art method for the generation of time series data. Deep generative models can learn underlying temporal and statistical correlations within the data and make use of substation metadata along with weather and calendar variables, enabling higher-quality load profile scenario generation. Secondly, this study considers the synthesis of reactive power profiles at the LV substation level. Reactive power is a key part of load behaviour and technologies such as heat pumps; the assumption of a flat power factor is not realistic, as specific appliances generate a reactive power requirement, the use of which is driven by behavioural routine. Coupling of routine behaviours means that this relation is not necessarily 1-to-1. Furthermore, the synthesis of reactive power facilitates a power systems analysis. Unlike in other similar works, the evaluation of realism is not entirely down to statistical measures – co-occurrence of load behaviours across cohorts of synthesised LV substation will affect the load and voltage profile at higher levels of the network, hence a power system evaluation needs to be carried out to ensure that diversity (and synchronisation of premises or substation extreme events) is realistic. The proposed model can capture and replicate realistic substation-specific behaviour without the significant overhead of mass data collection and storage.
@@ 3 Methodology
3.1 SSSDS4 Diffusion Model
Diffusion models were chosen as the generative approach for synthesis in this study owing to their excellent performance in imputation and generation across various data modalities. Furthermore, diffusion models have previously outperformed GANs and CVAEs specifically in time series data [26, 27, 28, 29, 30].
@@ 3.1 SSSDS4 Diffusion Model
The proposed diffusion model is built upon an implementation of the SSSDS4 model [26]. SSSDS4 provides a time series implementation of the diffusion model enhanced with the combination of structured state space models, to allow capturing long-term dependencies in the data, such as seasonal and diurnal trends linked by weather and behavioural routine [33]. The SSSDS4 implementation of diffusion models was chosen due to its state-of-the-art performance in several generation tasks across multiple datasets when benchmarked against other diffusion-based approaches [26]. The SSSDS4 model was adapted to utilise conditional inputs to constrain the generation process and synthesise samples from specific areas of the overall distribution. Conditions are introduced to the model as additional channel inputs and utilise the binary masking mechanism to discern signal and conditional inputs. When the mask value is 1, the original input is provided to the model as a condition; when it is 0, the input is replaced with noise, and the model must learn to denoise it. The model then learns a transformation from the conditional inputs and the signal noise vector to a novel load profile.
Furthermore, the SSSDS4 model source code contained an implementation error in the training loop relating to the handling of conditional inputs. Algorithm 1 outlines the modified training step in the SSSDS4 model. There is one key difference between this and the previous model. Line 7 restores the conditional values after the entire sample is corrupted with noise. Previously, this was handled differently; the conditional values were substituted into $x_{std}$ (the noise vector). Then, when the sample is corrupted, the conditional values would not be corrupted with noise. However, the noise corruption equation outlined in line 6 only balances the data for values of $x_{0}$ $\in[0,1]$. In many cases, this is not sufficient; a more effective method is to corrupt the entire sample with noise, then replace the values after corruption. This aligns with the behaviour during inference, where the conditional values are replaced at each step.
@@ 3.2 Load Profile Synthesis Models
Each model was trained with the optimal neural network configuration defined in [26], with 200 diffusion steps. Each model was trained with a maximum of 200 epochs, but was halted earlier if test loss plateaued. Figure 4 shows the convergence of each model based on the MSE vs the number of epochs; LVGenU converges relatively quickly because there are no conditions for the model to learn from; LVGenWC and LVGenWCS require a significantly higher number of iterations before the models converge.
@@ 3.3 Data Set Description
Distribution substation monitoring data was acquired from the National Grid Electricity Distribution (NGED) OpenData [34] platform. It constitutes a subset of monitored LV Networks consisting of 1,431 substations, which total to approximately 19,050 days of data (as of 6th of June 2024). Measurements include the active and reactive power for each phase (L1, L2, L3). Data for each substation was also paired with its associated metadata (such as primary substation number, number of connected customers, LCT capability, etc.), weather from the Meteostat API [35] and calendar information. Where there were missing or incomplete data samples, a discard strategy was adopted. Data cleansing also removed outliers such as data with unrealistically high values (greater than 1,000 kW), incomplete days, and other anomalous data (such as constant values recorded throughout the day), resulting in a cleansed dataset of 13,756 days. The dataset was split into training and testing sets based on the number of observations per substation, placing substations with a low number of observations in the testing set. This was to inhibit overfitting due to limited training data. A cut-off of 10 days was chosen to achieve a balanced split, with observations having fewer than 10 days placed in the testing set. This threshold was selected to maintain an approximate split of 70% for training and 30% for testing. Furthermore, this method of splitting was also chosen to ensure representative data synthesis at the substation level; data segments from the same substation could not be included in both the training and testing sets, i.e. data leakage was prevented by allocating all data from a particular substation to either the test or train sets.
The number of customers served by a substation was captured through fixed-width bins: 0-99, 100 - 199, etc.; the final bin contained all substations with greater than 600 customers. This form of discretisation was performed to prevent the model from memorising specific behaviours of individual substations based on their associated number of customers. This, in turn, enables the model to generalise and scale to potentially larger network scenarios. To create representative training/test sets, the split was performed in a stratified fashion based on the discrete group of the number of customers. Finally, in order to permit testing of the load flow simulation presented in Section 5, a set of substations which are connected to the same Primary Substation were placed in the testing set.
@@ 3.4 Evaluation Metrics and Visualisations
In this study, two metrics are used to evaluate the performance of the data synthesis models in two key areas. Mean Squared Error (MSE) is used to measure the reconstruction errors of the generated load profiles vs the actual. MSE was chosen as it is an error metric that is routinely used with machine learning applications. The Maximum Mean Discrepancy (MMD) is used to measure the quality of the synthesis. MMD is a distance-based metric measuring the difference between real and generated probability distributions by comparing their means in a high-dimensional feature space [36]. A lower score represents a more similar data distribution. MMD was chosen as it is regularly used in data synthesis studies as a distance-based metric [37]. Additional metrics, inspired by previous research [38, 39], were also considered and are listed in the table. These further distance-based metrics Wasserstein distance [40], marginal score and MiVo [38]. The remainder of this section focuses on MSE and MMD metrics, with the additional metrics referenced only when they provide additional insight.
@@
Data visualisations are used alongside these metrics to further evaluate the quality of the generated samples. The first kind of visualisation used is distribution plots, which contrast the implied probability distribution of the actual data and each proposed model. A zoomed-in section of the tail of the distribution is also provided to inspect the important extreme values, which in operational scenarios could result in problems such as voltage collapses. The second method of visualisation is demonstrated through decile plots, which plot the values of the load at various deciles of the implied probability distribution. The final visualisations used are Autocorrelation Function (ACF) plots, which measure the correlation of the data with a lagged version of itself across various time intervals. Similar ACF plots indicate that datasets share a similar underlying temporal structure, meaning they exhibit comparable diurnal patterns or intra-day dependencies. Thus, the resemblance in ACF plots implies that the datasets may have been generated by similar processes or are subject to similar temporal influences such as behavioural routine or local weather conditions [41, 19].

### v2

@@ Abstract
Limited visibility of distribution network power flows at the low voltage level presents challenges to both distribution network operators from a planning perspective and distribution system operators from a congestion management perspective. More representative loads are required to support meaningful analysis of LV substations; otherwise, such analysis risks misinforming future decisions. Traditional load profiling relies on typical profiles, oversimplifying substation-level complexity. Generative models have attempted to address this through synthesising representative loads from historical exemplars; however, while these approaches can approximate load shapes to a convincing degree of fidelity, analysis of the co-behaviour between substations is limited, which ultimately impacts higher voltage level network operation. This limitation will become even more pronounced with the increasing integration of low-carbon technologies, as estimates of base loads fail to capture load diversity. To address this gap, Conditional Diffusion models for synthesising daily active and reactive power profiles at the low voltage distribution substation level are proposed. The evaluation of fidelity is demonstrated through conventional metrics capturing temporal and statistical realism, as well as power flow modelling. Multiple models are proposed to handle varying levels of data availability, ranging from unconditional synthesis to an informed generation driven by metadata and daily statistics. The results show synthesised load profiles are plausible both independently and as a cohort in a wider power systems context. The Conditional Diffusion model is benchmarked against both naive and state-of-the-art models to demonstrate its effectiveness in producing realistic scenarios on which to base sub-regional power distribution network planning and operations.
@@ 1 Introduction
Scenario generation is one viable route to forestall potential problems from the introduction of new technology or general load growth. However, a baseline performance measure is difficult to obtain due to the lack of LV observability. At the distribution level, the primary concern is the extremes: headroom for thermal exceedances and footroom for voltage violations [3]. Typical Load Profiles (TLP) exist [10], but these are generally seasonal, not weather sensitive and lack the diversity that would enable a cohort of premises to be aggregated together into a plausible LV substation load. In defining plausibility, the peak demand sharpness (and under embedded PV scenarios, trough depth) should reflect the range of time use across neighbourhood premises: too sharp would result from near identical behaviours and hence an overestimate of potential threats; too blunt assumes greater diversity than is likely, and therefore a potential overestimate of headroom. Generative modelling from a limited set of historical exemplars offers a potential solution, but the varied nature of LV substation loading introduces its own set of complexities and hurdles. The statistical distribution of LV substation load does not follow any common probability density function [34], making sampling approaches unrepresentative unless carefully chosen for particular locations. Markov Chain approaches also struggle with capturing aggregated time use resulting from routine behaviours [21]. Understanding the present loading characteristics is essential as this condition could then further worsen with the extensive installation of distributed generators as well as low-carbon heat and transport [23]. This will cause equipment to operate under atypical conditions that impact the wider network [11], potentially up to the transmission level. For a Distribution Network Operator (DNO), this insight is necessary for long-term planning of reinforcements that will facilitate Low Carbon Technology (LCT) adoption and generally increase resilience over time periods of months, years and decades. For a Distribution System Operator (DSO), the same insights will yield short-term generation dispatch and load curtailment potential within the constraints of the physical network infrastructure, with both active and reactive turn-up and turn-down capabilities articulated accurately via a flexibility envelope  [33]. This, in turn, will articulate substation flexibilities that can provide balancing at the transmission interface, which, combined, represent regional-scale scenarios from fundamental end-use insights - unlocking this knowledge of aggregating demand would enable better system planning as well as policy design over a longer period.
@@
LV distribution network loads are much closer to individual premises end use, meaning that lower levels of aggregation will reflect highly variable behaviour routines which may result from social norms, variation in appliance ownership and specification, and also reactions to localised weather conditions. Previously, LV substation data was modelled using conventional methods such as aggregated average profiling [10] with a diversity correction. While this approach was historically sufficient, the increasing penetration of LCTs now necessitates more advanced and accurate methods. To approach this complex and non-stationary behaviour, it is proposed here that a generative AI model, specifically a diffusion-based model, is applied. This is capable of synthesising both Active and Reactive power data for LV substation load data based on a set of observed influencing factors, namely, weather and calendar variables. Diffusion is suitable for application on load data due to its ability to learn complex underlying temporal and statistical correlations present in load profile data. Furthermore, a conditional diffusion model can make use of substation metadata along with weather and calendar variables, enabling higher-quality load profile scenario generation. Diffusion is also the state-of-the-art method for the generation of time series data.
Multiple diffusion models are proposed throughout this study, each with their own set of conditional variables. This design choice is motivated by two main factors: variability in data availability across substations and the expected gains in generative fidelity offered by conditioning. Firstly, in reality, available data will vary across substations, ranging from limited monitoring to basic metadata or, in some cases, no information at all. This study aims to provide multiple models capable of handling all these scenarios, both for practical deployment reasons and to demonstrate the expected improvements in generative fidelity that arise from conditioning on additional information. The proposed model can capture and replicate realistic substation-specific behaviour by conditioning on lightweight information such as basic metadata or daily summary statistics, without the significant overhead of mass data collection and storage. These models will also be capable of supporting different use cases, such as scenario generation under limited data availability or unconditional and unrestricted data generation for complete hypothetical network generation, which will be outlined later in this study.
In this study, coherence is not defined as the joint synthesis of an entire LV network, nor as explicit coordination between substations during generation. Instead, it is treated as an emergent property arising from the aggregation of independently generated LV substation load profiles. The proposed diffusion models synthesise individual substation active and reactive power profiles with realistic temporal structure and diversity learned from historical data, without enforcing synchronisation.
Coherence is therefore evaluated a posteriori via power flow modelling, by propagating the independently synthesised substation loads through a sample distribution network and examining the resulting voltage magnitudes and phase angles. This analysis serves as a test of whether the generated loads collectively produce physically plausible and operationally meaningful network behaviour. Synchronisation would exaggerate voltage and thermal stress, while excessive smoothing would mask potential constraint violations. The objective is thus to generate substation-level loads whose diversity is sufficient to yield realistic network responses when analysed using standard power system tools.
1.1 Summary of Contributions
This section will outline the key contributions made in this study. These are as follows:
  1. 1.
    Application of a deep generative diffusion-based model to LV substation load modelling - To the best of the authors’ knowledge, this study represents the first application of diffusion-based generative modelling to load synthesis at the LV substation level, addressing a scale and use case not considered in prior diffusion-based load studies, which have primarily focused on other parts of the power system.
  2. 2.
    Synthesis of reactive power at the LV substation level - This study is the first to extend LV substation load synthesis to reactive power. Modelling reactive power is becoming increasingly important as it captures the aggregated inductive and capacitive characteristics of devices connected at the LV level and is therefore a critical quantity for realistic LV load modelling; as emerging technologies are increasingly connected to the grid (e.g., heat pumps, EVs, and other low‑carbon technologies), reactive power is expected to become more variable at the LV level, such that the previously common assumption of a constant power factor no longer holds.
  3. 3.
    Power system analysis of synthesised LV load profiles - Unlike many existing studies, realism is not assessed solely through statistical similarity of the individual load profiles. Instead, the proposed approach explicitly examines how the co-occurrence of load behaviours across cohorts of independently synthesised LV substations propagates through the network and affects voltage magnitude and phase angle behaviour at higher voltage levels using a sample network. This power-system based evaluation tests coherence and co-behaviour in the power systems context, which results in a more realistic operational consequence on which to base a planning or investment decision.
@@ 3 Methodology
3.1 SSSDS4 Conditional Diffusion Model
Diffusion models were selected as the deep learning generative framework because their modelling assumptions align well with the statistical and temporal characteristics of LV substation load data. LV load profiles are high-dimensional, strongly time-dependent, and exhibit non-Gaussian, often multi-modal distributions driven by aggregated behavioural routines and external conditions. Diffusion models have demonstrated strong performance in time-series generation and imputation tasks across multiple domains, consistently outperforming GANs and CVAEs in capturing complex temporal structure while preserving sample diversity [1, 36, 7, 8, 17]. These improvements were demonstrated across multiple datasets in different domains, showing the strong robustness and generalisation of the models. Diffusion models are particularly applicable to load data due to their ability to model high-dimensional data distributions while maintaining their diversity [43]. These properties are particularly important for LV load synthesis, where a realistic representation of diversity is required to avoid both excessive synchronisation and unrealistically smooth aggregation effects. Accordingly, diffusion models were chosen as a principled and suitable framework for the specific problem of LV substation load synthesis.
Conditional diffusion was adopted to exploit contextual information that is typically available in distribution network settings, such as calendar variables, weather data, and basic substation metadata, to guide the generation process. While diffusion models provide improved diversity relative to GAN-based approaches [43], accurately representing the heterogeneity and substation-specific behaviour of LV loads remains challenging in an unconditional setting. Introducing conditioning constraints the generative process to physically and contextually plausible regions of the data space, enabling more faithful representation of substation-level diversity and improved reconstruction accuracy. The effectiveness of conditional diffusion for load-related modelling tasks has also been demonstrated at other levels of the power system  [48] outside the LV level of aggregation.
@@ 3.1 SSSDS4 Conditional Diffusion Model
The proposed diffusion model is built upon an implementation of the SSSDS4 model [1]. SSSDS4 provides a time series implementation of a diffusion model and was shown to achieve strong performance in several generation tasks across multiple datasets when benchmarked against other diffusion-based approaches [1]. SSSDS4 was enhanced over existing methods [37] through the incorporation of structured state space models, to allow capturing long-term dependencies in the data, such as seasonal and diurnal trends linked by weather and behavioural routine [15]. These are known to be key factors in load behaviour [32] and were the key motivation in the selection of the SSSDS4 as the diffusion model to apply to the LV substation synthesis.
The SSSDS4 model was adapted to utilise conditional inputs to constrain the generation process and synthesise samples from specific areas of the overall distribution. Details of the masking mechanism are outlined in B. Furthermore, the SSSDS4 model source code contained an implementation error in the training loop relating to the handling of conditional inputs. Algorithm 1 outlines the modified training step in the SSSDS4 model. There is one key difference between this and the previous model. Line 7 restores the conditional values after the entire sample is corrupted with noise. Previously, this was handled differently; the conditional values were substituted into $x_{std}$ (the noise vector). Then, when the sample is corrupted, the conditional values would not be corrupted with noise. However, the noise corruption equation outlined in line 6 only balances the data for values of $x_{0}$ $\in[0,1]$. In many cases, this is not sufficient; a more effective method is to corrupt the entire sample with noise, then replace the values after corruption. This aligns with the behaviour during inference, where the conditional values are replaced at each step.
@@ 3.2 Load Profile Synthesis Models
Each model was trained with the optimal neural network configuration defined in [1], with 200 diffusion steps. Each model was trained with a maximum of 200 epochs, but was halted earlier if test loss plateaued. Figure 4 shows the convergence of each model based on the MSE vs the number of epochs; LVGenU converges relatively quickly as training terminates when no improvement in validation loss is observed for 2 consecutive checkpoints. A short patience was selected because the unconditional model rapidly reaches a performance plateau, after which further training does not yield measurable improvements and risks overfitting rather than learning additional structure. This occurs because there are no conditions for the model to learn from; LVGenWC and LVGenWCS require a significantly higher number of iterations before the models converge. Full details of the model configuration are available in C.
@@ 3.3 Data Set Description
Distribution substation monitoring data was acquired from the National Grid Electricity Distribution (NGED) OpenData [28] platform. It constitutes a subset of monitored LV Networks consisting of 1,431 substations, totalling approximately 19,050 days of data (as of 6th of June 2024). Measurements include the active and reactive power for each phase (L1, L2, L3). Each substation was paired with associated metadata, including primary substation identifier, number of connected customers, weather variables obtained from the Meteostat API [27], and calendar information. Records with missing or incomplete samples were discarded. Data cleansing further removed daily profile outliers (e.g., days with unrealistically high load values exceeding 1,000 kW), incomplete days, and anomalous profiles such as constant values recorded throughout the day. If any of these criteria were met, the entire daily load profile was removed. This resulted in a cleansed dataset of 13,756 days.
The dataset was split into training and test sets at the substation level. For each substation, the total number of available daily profiles was recorded, and substations with fewer than 10 days of data were assigned to the test set, while the remaining substations were assigned to the training set. This threshold was chosen to mitigate overfitting/memorisation during training. Using a cut-off of 10 days resulted in an approximate 70/30% train–test split (69.93% of samples in the training set). This strategy also prevents data leakage, as all observations from a given substation are contained exclusively within a single split.
The number of connected customers was discretised into fixed-width bins (0–99, 100–199, etc.), with a final bin aggregating substations with more than 600 customers. This discretisation was adopted to prevent the model from memorising or implicitly regressing on the exact customer count, which would reduce generalisation and risk information leakage across substations with similar sizes. This upper bin was introduced to avoid sparsely populated bins at the upper end of the customer-count distribution, as fewer than 20 substations exceeded this threshold (representing approximately 0.83% of all substations, with a maximum of approximately 800 customers).
Finally, in order to permit testing of the load flow simulation presented in Section 5, a set of substations which are connected to the same Primary Substation were held back from the testing set.
@@ 3.4 Evaluation Metrics and Visualisations
In this study, two metrics are used to evaluate the performance of the data synthesis models in two key areas. Mean Squared Error (MSE) is used to measure the reconstruction errors of the generated load profiles vs the actual. MSE was chosen as it is an error metric that is routinely used with machine learning applications. The Maximum Mean Discrepancy (MMD) is used to measure the quality of the synthesis. MMD is a non-parametric, kernel-based distance metric that quantifies differences between real and generated probability distributions by comparing their means in a reproducing kernel Hilbert space [14]. LV substation load data are non-Gaussian and do not follow a common parametric probability density function [34], rendering parametric divergences (e.g., KL divergence) unsuitable or unstable in this context. MMD was therefore selected as it enables robust comparison of complex, multi-modal load distributions without imposing restrictive distributional assumptions. For this reason, MMD is widely adopted in data synthesis studies involving non-Gaussian and high-dimensional data [12].
The remainder of this section will focus on MSE and MMD metrics; however, additional metrics are also reported, mainly inspired by previous research in load synthesis, to provide further insights into model performance [2, 24]. These further distance-based metrics include Wasserstein distance [41], which can further evaluate modelling at the distribution tail, marginal score [24], which models the distributional alignment based on a fixed number of bins, and MiVo [2], which provides a distance-based metric with additional awareness to diversity. These metrics can be used alongside the core metrics to better understand the modelling of key areas of the distribution and will be referenced where appropriate. An additional tail-aware MSE metric is proposed to quantify discrepancies between generated and real data at distributional extremes. This is computed by restricting the MSE calculation to observations exceeding the 95th percentile of the empirical distribution. This metric provides targeted insight into the modelling of extreme operating regions, which are particularly relevant in load modelling to avoid issues such as thermal exceedances and voltage collapse. This metric is referred to as the Tail-Focused MSE.
Each metric is computed on the full test set for both active and reactive power, with the average value reported as the final score. The MSE, Wasserstein distance, and Marginal Score are computed on daily aggregated distributions collated across all test-set substations, providing a global assessment of synthesis quality. In contrast, the MMD and MiVo metrics are computed on daily profiles of individual substations for the joint distribution of active and reactive power, with the reported value corresponding to the average across all substations. The MSE and Wasserstein distances are computed using standard Python implementations from scikit-learn (v0.22.2) and SciPy (v1.16.3), respectively. For MMD and Marginal Score, the methodology described in [24] is adopted, while MiVo is computed following the approach outlined in [2]. Separate performance of active and reactive power on each metric is available in Appendix D.
@@
Data visualisations are used alongside these metrics to further evaluate the quality of the generated samples. The first kind of visualisation used is distribution plots, which contrast the implied probability distribution of the actual data and each proposed model. A zoomed-in section of the tail of the distribution is also provided to inspect the important extreme values, which in operational scenarios could result in problems such as voltage collapses. The second method of visualisation is demonstrated through decile plots, which plot the values of the load at various deciles of the implied probability distribution. The decile plots (10th–90th percentiles) are calculated across substations by extracting the corresponding decile at each timestep. This produces representative load profiles for each decile, illustrating variability and tail behaviour over the diurnal cycle. The final visualisations used are Autocorrelation Function (ACF) plots, which measure similarities in temporal dependence between real and generated datasets. The ACF is computed independently for each substation load profile by correlating the signal with lagged versions of itself. This is performed at each 10-minute interval of the daily load profile, resulting in 143 lag values per substation. The ACF values are then averaged across substations to provide an aggregate comparison of temporal dependency structure between models. Similar ACF plots indicate that datasets share a similar underlying temporal structure, meaning they exhibit comparable diurnal patterns or intra-day dependencies. Thus, the resemblance in ACF plots indicates that the generated datasets preserve similar temporal dependency patterns to those observed in the real data [4, 44].